OpenStudy (sleepyjess):

Proving trig equations (again) \(\sf cotx~sec^4x=cotx+2tanx+tan^3x\)

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (mathmath333):

lol it was faulty \(\large\tt \begin{align} \color{black}{ = \cot x\sec^4x\\~\\ =\cot x(\sec^2x)^2\\~\\ =\cot x(1+\tan^2x)^2\\~\\ =\cot x(1+2\tan^2x+\tan^4x)\\~\\ }\end{align}\)

OpenStudy (michele_laino):

If I multiply both sides by tan x, I get: \[\frac{ 1 }{ (\cos x)^{4} }=1+2(\tan x)^{2}+(\tan x)^{4}=[1+(\tan x )^{2}]^{2}\]

OpenStudy (anonymous):

Yeah, there we go @mathmath333

OpenStudy (mathmath333):

yea its ok , now i think

OpenStudy (sleepyjess):

I see where that is coming from, but how would that become \(\sf cotx+2tanx+tan^3x\)?

OpenStudy (mathmath333):

further u have to multiply \(\cot x\) in the bracket and use this \(\cot x=\dfrac{1}{\tan x}\)